OpenStudy (anonymous):

What is the value of the expression? 4^3 • 4^−5 A. 4^−15 B. 4^−2 C. 4^2 D. 4^8

OpenStudy (anonymous):

@gerryliyana

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (ribhu):

a^n . a^m = a^m+n

OpenStudy (anonymous):

No not really.

OpenStudy (ribhu):

yeah it is.

OpenStudy (anonymous):

Im talking about any ideas.

OpenStudy (skullpatrol):

$$\Huge4^3 • 4^{−5}$$

OpenStudy (anonymous):

4^3 . 4^−5 = (4x4x4)x(1/(4x4x4x4x4))

OpenStudy (ribhu):

4^-2

OpenStudy (anonymous):

I'm so confused, this is what i got 0.0625

OpenStudy (anonymous):

\[4^3 \times 4^5 = (4 \times 4 \times 4) \times \frac{ 1 }{ 4 \times 4 \times 4 \times 4 \times 4 }\]

OpenStudy (anonymous):

I have a 5 question test i need to get this done asap

OpenStudy (skullpatrol):

$$\Huge\frac{4^3}{ 4^5}$$

OpenStudy (anonymous):

Im failing in my grade i need to get this done

OpenStudy (anonymous):

\[4^3 \times 4^{-5} = \frac{ 1 }{ 4 \times 4 \times 4 \times 4 \times 4 }\]

OpenStudy (ribhu):

@Kyliejones i wrote an answer its B

OpenStudy (anonymous):

lot of verbiage, but the bottom line is you add the exponents

OpenStudy (anonymous):

ahhhh latex is suck!

OpenStudy (anonymous):

Thank you all of you im failing and i need to get this done asap

OpenStudy (anonymous):

\[4^3 \times 4^{-5} = (4 \times 4 \times 4) \times \frac{ 1 }{ 4 \times 4 \times 4 \times 4 \times 4 }\]

OpenStudy (anonymous):

the answer is 4^-2

OpenStudy (anonymous):

it's B!!!